# Solving and learning advective multiscale Darcian dynamics with the Neural Basis Method

### v1

@@ Methods
$\|u-v_{N_{b}}\|_{V}\;\leq\;C_{\mathrm{emb}}\,\|u-v_{N_{b}}\|_{\mathcal{H}_{k}}.$ (M8)
Combining (M7) and (M8) yields
$\|u-v_{N_{b}}\|_{V}\;\leq\;C_{\mathrm{emb}}\,C_{1}(\delta)\,\|u\|_{\mathcal{H}_{k}}\,N_{b}^{-1/2}.$ (M9)
@@
$\inf_{w\in V_{N_{b}}}\|u-w\|_{V}\;\leq\;\|u-v_{N_{b}}\|_{V}.$ (M10)
Substituting (M10) into the Céa’s estimate (M6) and using (M9) gives
@@
$\|u-u_{N_{b}}\|_{V}\;\leq\;\frac{C_{a}}{\gamma}\left(\inf_{w\in\mathcal{H}_{k}}\|u-w\|_{V}+C_{\mathrm{emb}}\,C_{1}(\delta)\,N_{b}^{-1/2}\right),$
which completes the proof. ∎
@@
Collocation implementation of mixed weighted least-squares for multiscale Darcy flow
Each residual evaluation at a collocation point corresponds to one row of the discrete weighted least-squares system (5). For interior points $x_{i}\in\Omega$ with associated volume $|\Omega_{i}|$, we use $w_{m}(x_{j})=1/\rho\,\sqrt{\mu/\kappa(x_{j})}\,\sqrt{|\Omega_{j}|}$ for the constitutive residual and $w_{c}(x_{j})=\ell/\rho\,\sqrt{\mu/\kappa_{\mathrm{mean}}}\,\sqrt{|\Omega_{j}|}$ for the continuity residual. For boundary points $x_{b}\in\partial\Omega$ with associated length or area $|\partial\Omega_{b}|$, we set $w_{D}(x_{b})=\sqrt{1/h}\,\sqrt{\kappa_{\mathrm{mean}}/\mu}\,\sqrt{|\partial\Omega_{b}|}$ and $w_{N}(x_{b})=\ell/\rho_{ref}\,\sqrt{1/h}\,\sqrt{\mu/\kappa_{\mathrm{ref}}}\,\sqrt{|\partial\Omega_{b}|}.$ For collocation points that are uniformly distributed over $\Omega$ and $\partial\Omega$, the measures $|\Omega_{i}|$ and $|\partial\Omega_{b}|$ are straightforward to compute. Accordingly, the discrete residual vector takes the stacked form
@@
where each row of $\mathbf{A}_{m},\mathbf{A}_{c},\mathbf{A}_{D},\mathbf{A}_{N}$ is obtained by evaluating the bases and their derivatives at the corresponding collocation point and inserting them into the constitutive, continuity, and boundary residual definitions. Finally, the weighting matrix in (5) is diagonal under collocation
@@
where $\mathbf{w}_{m}$ collects $w_{m}(x_{j})$ on interior points, $\mathbf{w}_{c}$ collects $w_{c}(x_{j})$ on interior points, and $\mathbf{w}_{D}$ and $\mathbf{w}_{N}$ collect $w_{D}(x_{b})$ and $w_{N}(x_{b})$ on Dirichlet and Neumann boundary points.
Supercritical CO2 storage models
We consider a supercritical CO2 storage benchmark in a square porous domain $\Omega=[0,762]~\mathrm{m}\times[0,762]~\mathrm{m}$ (Fig. 2a) under two permeability settings: (i) a homogeneous case with constant permeability $\kappa=0.3~\mathrm{D}$, and (ii) a heterogeneous case in which $\kappa(\mathbf{x})$ is generated as a geologically realistic Gaussian random field. The heterogeneous permeability field is shown in Fig. 3a and has a permeability contrast of $502$. A passive tracer is injected at the left boundary. Except for the permeability distribution, the rest of the setups of the two cases are identical.
Boundary conditions: on the left boundary ($x=0$), we set constant pressure $p=103.4~\mathrm{bar}$ and impose a passive tracer pulse with concentration $c_{\mathrm{in}}=100~\mathrm{ppm}$ lasted for the first 10 days. On the top boundary ($y=762~\mathrm{m}$), we impose an outward mass-flux $q_{y}=586~\mathrm{kg~day^{-1}~m^{-2}}$. The right boundary ($x=762~\mathrm{m}$) is no-flow ($q_{x}=0$), and the bottom boundary ($y=0$) is no-flow ($q_{y}=0$).
Fluid and rock properties: rock and fluid properties are uniform unless otherwise stated by the permeability setting: porosity $\varepsilon=0.25$, fluid viscosity $\mu=0.04~\mathrm{cP}$, compressibility $c_{f}=2.90\times 10^{-3}~\mathrm{bar}^{-1}$, and reference density $\rho_{0}=384~\mathrm{kg\,m^{-3}}$ at $50~^{\circ}\mathrm{C}$ and $p_{0}=103.4~\mathrm{bar}$.
Reference and discretizations: for the homogeneous case, we use an FVM solution computed on a $1000\times 1000$ structured uniform mesh as the ground-truth, which provides a reliable numerical proxy for the true solution in this setting. For the heterogeneous case, a ground truth analogous to the homogeneous fine-grid proxy is not available by grid refinement, because the permeability is discretely-specified through a generated heterogeneous realization rather than a known true continuous field. For both the homogeneous and heterogeneous cases, the baseline FVM solutions uses a $50\times 50$ structured uniform mesh. For NBM, we employ collocation on a $50\times 50$ uniform interior grid (2500 interior points) to match the spatial resolution of the $50\times 50$ FVM baseline discretization. We place $50$ uniformly distributed collocation points on each boundary, yielding $2700$ total collocation points. Vanilla PINN baselines using the same spatial resolution are provided in SI.8. For the comparison between using and not using the energy-consistent weight (4), the baseline is uniform weighting with constant weights of 1.
Manufactured-solution benchmark with separable multiscale permeability
To quantify intrinsic accuracy in a multiscale setting where grid refinement does not provide a ground-truth proxy, we introduce a manufactured-solution benchmark. We consider a simplified incompressible case with a spatially multiscale yet separable permeability field. We consider
$\mathbf{u}=-\frac{\kappa(x,y)}{\mu}\nabla p,\qquad\nabla\cdot\mathbf{u}=0,\qquad(x,y)\in\Omega,$
posed on a regular domain $\Omega$ with constant viscosity $\mu$. The permeability is constructed as a separable product $\kappa(x,y)=a(x)\,b(y)$, where $a$ and $b$ are oscillatory 1D fields to induce controlled heterogeneity and contrast while remaining smooth within each coordinate. Specifically, we use,
$a(x)=\exp\!\left(s_{x}\frac{\varphi_{x}(x)}{\max|\varphi_{x}|}\right),\qquad b(y)=\exp\!\left(s_{y}\frac{\varphi_{y}(y)}{\max|\varphi_{y}|}\right),$
with $\varphi_{x}(x)=\beta_{x1}\sin(2\pi x)+\beta_{x2}\cos(4\pi x)$ and $\varphi_{y}(y)=\beta_{y1}\cos(2\pi y)+\beta_{y2}\sin(2\pi y)$. The scaling parameters $s_{x}=\tfrac{1}{2}\log(\mathcal{C}_{x})$ and $s_{y}=\tfrac{1}{2}\log(\mathcal{C}_{y})$ control the 1D contrasts $\mathcal{C}_{x}$ and $\mathcal{C}_{y}$ and thereby the overall contrast. All permeability values are evaluated on cell centers of a uniform grid. We thus can manufacture an exact pressure field by defining 1D antiderivatives of the reciprocal coefficients,
$I_{x}(x)=\int_{x_{\min}}^{x}\frac{1}{a(s)}\,ds,\qquad I_{y}(y)=\int_{y_{\min}}^{y}\frac{1}{b(s)}\,ds,$
and setting $p(x,y)=p_{0}+B\bigl(C_{x}I_{x}(x)+C_{y}I_{y}(y)\bigr)$, where $p_{0}$ is a reference pressure, $B$ sets the amplitude, and $C_{x},C_{y}$ specify the relative strengths of the $x$ and $y$ components. We thus have
$\partial_{x}p=B\,\frac{C_{x}}{a(x)},\qquad\partial_{y}p=B\,\frac{C_{y}}{b(y)},$
which implies an analytic velocity field
$u_{x}(x,y)=-\frac{\kappa(x,y)}{\mu}\partial_{x}p=-\frac{B\,C_{x}}{\mu}\,b(y),\qquad u_{y}(x,y)=-\frac{\kappa(x,y)}{\mu}\partial_{y}p=-\frac{B\,C_{y}}{\mu}\,a(x).$
Since $u_{x}$ depends only on $y$ and $u_{y}$ depends only on $x$, the divergence-free constraint holds identically. The corresponding boundary conditions for $p$ are taken from the manufactured pressure, and the benchmark admits an exact solution pair $(p,\mathbf{u})$ for any chosen contrasts and coefficients. For implementation, the permeability factors $a(x)$ and $b(y)$ are evaluated on cell-center coordinates, giving $\kappa_{ij}=k_{0}\,a(x_{i})b(y_{j})$ with a uniform scaling factor $k_{0}$. The integrals $I_{x}$ and $I_{y}$ are computed on the 1D grids to assemble $p_{ij}$ on cell centers, using either a cumulative trapezoidal rule on the uniform grid or a higher-accuracy piecewise Gauss–Legendre 8-point quadrature on each cell interval. The exact velocity is evaluated from the closed-form expressions implied by the manufactured construction, avoiding numerical differencing.
In the experiments of Fig. 3k–s, we set $\Omega=[-1,1]\times[-1,1]$ with $p_{0}=34~\mathrm{bar}$, $\mu=1.0~\mathrm{cp}$, B=20B=-20, and $(C_{x},C_{y})=(1.0,-0.6)$, and use $k_{0}=0.1$. The permeability contrasts of the 1D factors are prescribed as $\mathcal{C}_{x}=\max(a)/\min(a)=200$ and $\mathcal{C}_{y}=\max(b)/\min(b)=20$ under global normalization and mean preservation, yielding an overall contrast is about 1078, while varying $(\mathcal{C}_{x},\mathcal{C}_{y})$ generates permeability fields with different contrast levels. This manufactured benchmark only provides exact pressure and velocity. The concentration grounth-truth is obtained by solving the transport equation with an FVM discretization using first-order upwinding driven by the manufactured velocity field, since an exact closed-form concentration solution is generally difficult to construct for heterogeneous advection under general initial and boundary conditions.
Residual–error certificate for the weighted Darcy projection

### v2

@@ 2.1 Core NBM formulation
$\|u-v_{N_{b}}\|_{V}\;\leq\;C_{\mathrm{emb}}\,\|u-v_{N_{b}}\|_{\mathcal{H}_{k}}.$ (4)
Combining (3) and (4) yields
$\|u-v_{N_{b}}\|_{V}\;\leq\;C_{\mathrm{emb}}\,C_{1}(\delta)\,\|u\|_{\mathcal{H}_{k}}\,N_{b}^{-1/2}.$ (5)
@@
$\inf_{w\in V_{N_{b}}}\|u-w\|_{V}\;\leq\;\|u-v_{N_{b}}\|_{V}.$ (6)
Substituting (6) into the Céa’s estimate (2) and using (5) gives
@@
2.2 Mixed least-squares with energy-consistent weights for multiscale Darcy flow
Discrete neural gradient and divergence operators are generally not adjoint. Consequently, a pressure-only approximation that eliminates $\mathbf{q}$ through (A.1) and then recovers it a posteriori does not, in general, preserve local mass conservation when inserted into (A.2). We therefore devise a mixed formulation and solve $\boldsymbol{\theta}_{p}$ and $\boldsymbol{\theta}_{q}$ simultaneously using least-squares for each update by
Here $\boldsymbol{\theta}_{p,q}$ denotes $(\boldsymbol{\theta}_{p},\boldsymbol{\theta}_{q})$, $p_{NN}=\sum_{i}\theta_{p,i}\,\phi_{p,i}$ and $\mathbf{q}_{NN}=\sum_{i}\theta_{q,i}\,\phi_{q,i}$, a compact form of (1). This mixed projection jointly enforces the constitutive law, the continuity constraint, and the boundary conditions, and produces locally conservative fluxes without a posterior correction. Additionally, $\|\cdot\|_{\Omega/\partial\Omega}$ denotes a generic norm defined in $\Omega$ or on $\partial\Omega$, and its construction hinges on how we measure the interior and boundary residuals.
Unweighted $L^{2}$ norms do not respect the Darcian energy scaling, which can degrade accuracy in multiscale media; this motivates energy-consistent weighting. Accordingly, we place all residual terms on a common Darcian energy scale: we weight the constitutive residual to match the dissipative metric $\int_{\Omega}\frac{\mu}{\kappa}\,|\mathbf{u}|^{2}\,d\Omega$, and scale the continuity and boundary residuals to the same metric. Under collocation realization, this yields:
with
In $w_{c}$ and $w_{N}$, $\ell$ is a characteristic length that ensures dimensional consistency relative to the constitutive residual. The weights $w_{D}$ and $w_{N}$ balance the pressure and mass-flux boundary residuals against the interior energy scaling with a measure correction $\sqrt{1/h}$ to compensate for the codimension-one boundary sampling in collocation. Here $h$ is the collocation spacing of the residual evaluation over the physical domain. $\kappa_{\mathrm{mean}}$ is the mean permeability and $\rho_{0}$ is the reference density. These weights measure all terms in compatible weighted $L^{2}$ norms and yield a consistent, physically interpretable projection.
In implementation, each residual evaluation at a collocation point corresponds to one row of the discrete weighted least-squares system. For interior points $x_{i}\in\Omega$ with associated volume $|\Omega_{i}|$, we use $w_{m}(x_{j})=1/\rho\,\sqrt{\mu/\kappa(x_{j})}\,\sqrt{|\Omega_{j}|}$ for the constitutive residual and $w_{c}(x_{j})=\ell/\rho\,\sqrt{\mu/\kappa_{\mathrm{mean}}}\,\sqrt{|\Omega_{j}|}$ for the continuity residual. For boundary points $x_{b}\in\partial\Omega$ with associated length or area $|\partial\Omega_{b}|$, we set $w_{D}(x_{b})=\sqrt{1/h}\,\sqrt{\kappa_{\mathrm{mean}}/\mu}\,\sqrt{|\partial\Omega_{b}|}$ and $w_{N}(x_{b})=\ell/\rho_{ref}\,\sqrt{1/h}\,\sqrt{\mu/\kappa_{\mathrm{ref}}}\,\sqrt{|\partial\Omega_{b}|}.$ Accordingly, the discrete residual vector takes the stacked form
@@ 2.2 Mixed least-squares with energy-consistent weights for multiscale Darcy flow
where each row of $\mathbf{A}_{m},\mathbf{A}_{c},\mathbf{A}_{D},\mathbf{A}_{N}$ is obtained by evaluating the bases and their derivatives at the corresponding collocation point and inserting them into the constitutive, continuity, and boundary residual definitions. Finally, we have the following diagonal weighting matrix
@@
where $\mathbf{w}_{m}$ and $\mathbf{w}_{c}$ collect the interior weights $w_{m}(x_{j})$ and $w_{c}(x_{j})$, while $\mathbf{w}_{D}$ and $\mathbf{w}_{N}$ collect the Dirichlet and Neumann boundary weights $w_{D}(x_{b})$ and $w_{N}(x_{b})$. We therefore write
$\boldsymbol{\theta}_{p,q}^{*}=\arg\min_{\boldsymbol{\theta}_{p,q}}\left\|\mathbf{W}\bigl(\mathbf{A}_{\mathrm{flow}}\,\boldsymbol{\theta}_{p,q}-\mathbf{b}_{\mathrm{flow}}\bigr)\right\|_{2}^{2}.$ (10)
We then define a residual measure in a relative sense as $\mathcal{E}_{\mathrm{rel}}=\big\|\mathbf{W}\bigl(\mathbf{A}_{\mathrm{flow}}\,\boldsymbol{\theta}_{p,q}-\mathbf{b}_{\mathrm{flow}}\bigr)\big\|_{2}^{2}/{\big\|\mathbf{b}_{\mathrm{flow}}\big\|_{2}^{2}}$.
2.3 Upwind control-volume stabilization for transport
For transport, we treat the velocity field $\mathbf{u}_{NN}$ as known and hold it frozen over each Darcy-flow update. To stabilize advection, we enforce (A.3) using a control-volume collocation with first-order upwinding. This gives the following discrete balance for each control-volume $K$
2.4 Implicit time integration and nonlinearity treatment.
For Darcy flow, given $p_{NN}^{n}$ at time $T^{n}$, we advance one Darcy time step $\Delta T$ to $T^{n+1}$ by solving (10) with implicit time integration. For the nonlinearity induced by pressure-dependent density in (A.1), we choose Picard linearization. At Picard iterate $k$, we update the coefficients by
3. Performance of NBM on single-query problems
In this section, we assess the performance of NBM on representative single-query Darcy–transport problems. We focus on solution accuracy, residual decay, basis refinement behavior, and robustness under increasing multiscale complexity. We begin with a homogeneous compressible benchmark to illustrate the basic approximation and convergence behavior of NBM, and then move to challenging multiscale permeability fields to examine the role of energy-consistent weighting and the robustness of the method in heterogeneous regimes.
3.1 NBM produces advective Darcian solutions with spectral-like accuracy
For a homogeneous compressible Darcy flow–transport problem based on supercritical CO2 storage (Fig. 2a), the residual $\mathcal{E}_{\mathrm{rel}}$ decays in two near-linear stages on semi-log scale as the neural basis size $N_{b}$ increases: a rapid drop that resolves dominant structures, followed by a milder slope as higher-frequency features become harder to represent with a finite basis (Fig. 2b). Consistent with this behavior, the relative $L_{2}$ errors of pressure, velocity, and concentration decrease rapidly and remain competitive with FVM across all variables (Fig. 2c). The velocity statistics also converge in a spectral-like manner, as quantified by the Kolmogorov–Smirnov distance, and the qualitative fields show systematic refinement with increasing $N_{b}$ (Fig. 2d–e). For fieldwise errors in pressure, velocity, and concentration, we compare NBM against a finite-volume method (FVM) and a vanilla PINN. NBM attains a $0.001\%$ relative $L_{2}$ error in pressure, about $50\%$ lower than FVM ($0.002\%$). For velocity, NBM yields $0.16\%$ for $u_{x}$ (FVM: $0.14\%$) and $0.30\%$ for $u_{y}$ (FVM: $0.33\%$), whereas the vanilla PINN underperforms ($2.65\%$ for $u_{x}$ and $4.43\%$ for $u_{y}$). For concentration, NBM ($5.17\%$) closely matches FVM ($5.36\%$), while the PINN degrades severely ($100\%$) (Fig. 2f–i). The experiment details of this example are provided in Appendix F.
Fig. 2: NBM performance on a homogeneous compressible benchmark.  a, Homogeneous CO2 storage model. b, Residual $\mathcal{E}_{\mathrm{rel}}$ vs. neural basis size $N_{b}$. c, Relative $L_{2}$ errors vs. $N_{b}$. d, Kolmogorov–Smirnov distance vs. $N_{b}$. e, Refinement of pressure and concentration solutions with $N_{b}$. f, Left to right: ground-truth pressure, followed by absolute error fields (w.r.t. ground truth) for NBM, FVM, and vanilla PINN. g, Left to right: ground-truth velocity magnitude, followed by absolute error fields (w.r.t. ground truth) for NBM, FVM, and vanilla PINN. h, Left to right: ground-truth concentration, followed by absolute error fields (w.r.t. ground truth) for NBM, FVM, and vanilla PINN. i, Relative $L_{2}$ errors. Note: All results are at 90 days. In f–i, NBM uses $N_{b}=1000$, and all methods (NBM, FVM, and vanilla PINN) employ a $50\times 50$ spatial resolution. Pressure is reported in $\mathrm{bar}$, velocity in $\mathrm{m\,day^{-1}}$, and concentration in $\mathrm{ppm}$.
Fig. 3: NBM performance on multiscale benchmarks.  a, Multiscale CO2 storage model. (permeability contrast $502$).  b, Residual $\mathcal{E}_{\mathrm{rel}}$ vs. neural basis size $N_{b}$.  c, Relative $L_{2}$ difference vs. $N_{b}$.  d, K-S distance vs. $N_{b}$.  e, Refinement of NBM pressure and concentration fields with increasing $N_{b}$.  f, Left to right: baseline pressure, velocity magnitude, and concentration.  g, Left to right: absolute differences of NBM with energy-consistent weighting, for pressure, velocity magnitude, and concentration.  h, Left to right: absolute differences of NBM without energy-consistent weighting, for pressure, velocity magnitude, and concentration.  i, Left to right: absolute differences of PINN, for pressure, velocity magnitude, and concentration.s.  j, Summary metrics measured against the baseline.  k, Manufactured-solution setup (permeability contrast 1078).  l, Residual $\mathcal{E}_{\mathrm{rel}}$ vs. $N_{b}$.  m, Relative $L_{2}$ error vs. $N_{b}$.  n, Energy spectrum $E(k)$ comparison.  o, Ground-truth pressure, velocity magnitude, and concentration fields.  p, Left to right: absolute errors of NBM, for pressure, velocity magnitude, concentration.  q, Left to right: absolute errors of FVM, for pressure, velocity magnitude, concentration.  r, Relative $L_{2}$ errors.  s, Comparing metrics vs. permeability contrast. Note: Results in b–j are at 90 days. In f–j and o-s, NBM uses $N_{b}=1000$, and all methods (NBM, FVM, and vanilla PINN) employ a $50\times 50$ spatial resolution. Pressure is reported in $\mathrm{bar}$, velocity in $\mathrm{m\,day^{-1}}$, and concentration in $\mathrm{ppm}$.
3.2 NBM remains accurate and robust for multiscale permeability fields
For this Darcy–transport problem but under a challenging multiscale permeability field (Fig. 3a), increasing the neural basis size $N_{b}$ yields similar spectral-like residual decay together with systematic basis refinement (Fig. 3b–e). Using the FVM solution as a baseline, fieldwise comparisons show that energy-consistent weighting is critical in the multiscale regime: without weighting, the flow-field energy-spectrum discrepancy $E(k)$ increases from $0.91\%$ to $7.05\%$ ($7.7\times$ increase), and the transport solution driven by this flow field deteriorates sharply, with the concentration relative $L_{2}$ difference rising from $4.09\%$ to $17.6\%$ ($4.3\times$) (Fig. 3f–j). By contrast, a vanilla PINN fails to produce meaningful solutions. Because a ground-truth is unavailable for this discrete-valued permeability field, the FVM solution is used only as a comparison baseline; discrepancies relative to it should not be interpreted as inaccuracy of NBM. We therefore include a manufactured-solution benchmark with a separable multiscale permeability to assess multiscale accuracy and robustness. NBM exhibits the same decay trend while converging to more accurate solutions than FVM (Fig. 3l–m), and the energy spectra confirm that it matches the dominant low-frequency mode of the true spectrum with much smaller spectral error ($0.31\%$ vs $8.95\%$) (Fig. 3n). NBM’s accuracy is also reflected in fieldwise errors: the relative $L_{2}$ errors drop from $0.078\%/4.02\%/6.74\%/4.52\%$ (FVM) to $0.052\%/0.61\%/1.03\%/0.51\%$ for $(p,u_{x},u_{y},c)$ (Fig. 3o–r). Moreover, robustness to increasing permeability contrast is demonstrated up to $10^{4}$, with variable-wise and spectral relative $L_{2}$ errors and the K–S distance remaining controlled across the tested contrasts (Fig. 3s). The experiment details of this example are provided in Appendices F, G and I.
4. NBM operator learning for parametric Darcy flow–transport
The results above establish that NBM is accurate and robust, but single-instance NBM solver requires dense least-squares solutions. With $N_{b}$ bases and $M$ residual evaluations, the computational cost scales as $\mathcal{O}(MN_{b}^{2})$, so NBM is not intended to compete with classical solvers in single-query settings. The situation changes in many-query regimes with parametric variations in PDE coefficients and boundary conditions, where repeated linear solves become the bottleneck in engineering design, control, and uncertainty analysis. NBM–OL addresses this by learning, offline, the parametric dependence of the solution coefficients in the predefined neural basis space and then evaluating them online at negligible cost.
We use NBM–OL to learn trajectory operators for the parametric system (A.5). Parametric encodings, together with time when relevant, are mapped to the neural basis coefficients, as summarized in Fig. 4. The parametric inputs include permeabilities $\boldsymbol{\xi}_{\kappa}$ and boundary values $\boldsymbol{\xi}_{b}$. To regularize learning and reduce dimensionality, we represent these inputs with low-dimensional encodings, as is standard in parametric PDE modeling [39]; for permeability, we use coarse blockwise parameters commonly adopted in geostatistical studies [23, 36]. The learned map predicts the neural basis coefficients, interpreted as the reduced coordinates of the solution in the NBM space. If additional compression is beneficial, we can apply proper orthogonal decompositio (POD) to the solution coefficient trajectories to further reduce the operator output dimension. The training process can be fully self-supervised, requiring no training data and thus avoiding dependence on expensive simulation databases. In addition, it minimizes the same residuals (13) and (14) used by the single-instance NBM solver for flow and transport. Notably, each residual provides an interpretable certificate for monitoring, diagnosis, and improving the learning process. Next, we first establish the respective equivalence relations with the solution errors using Lemmas 2 and 3 and then show that NBM–OL breaks the fundamental bottleneck of repeated parametric solves in many-query settings.
Fig. 4: NBM-OL for parametric Darcy–transport systems.  Panel I, Parameterization of permeability and boundary value. Panel II, NBM–OL for Darcy flow (per Darcy time step, $\Delta T$). $\mathrm{MLP}_{p}$ and $\mathrm{MLP}_{q}$ learn the pressure and mass-flux solution coefficients $\boldsymbol{\theta}_{\mathrm{p}}$ and $\boldsymbol{\theta}_{\mathrm{q}}$. Training is self-supervised by $\min\lVert\boldsymbol{R}(\boldsymbol{\theta}_{\mathrm{p}},\boldsymbol{\theta}_{\mathrm{q}})\rVert_{2}^{2}$ from (13), where $\boldsymbol{R}(\theta_{\mathrm{p}},\theta_{\mathrm{q}})=\mathbf{W}\bigl(\mathbf{A}_{\mathrm{flow}}\,\boldsymbol{\theta}_{p,q}-\mathbf{b}_{\mathrm{flow}}\bigr)$ with time stepping indices omitted. $p^{n}$ denotes the pressure from the previous Darcy time-step, used for the current update. Panel III, NBM–OL for transport (per transport time-step, $\Delta t$). $\mathrm{MLP}_{c}$ learns concentration coefficients $\boldsymbol{\theta}_{\mathrm{c}}$ by $\min\lVert\boldsymbol{R}(\boldsymbol{\theta}_{\mathrm{c}})\rVert_{2}^{2}$ from (14), where $\boldsymbol{R}(\theta_{\mathrm{c}})=\mathbf{A}_{\mathrm{up}}\,\boldsymbol{\theta}_{c}-\mathbf{b}_{\mathrm{up}}$ with time stepping indices omitted. $\mathbf{u}^{n+1}$ denotes the velocity after the current Darcy time-step update, computed as $\mathbf{u}^{n+1}=\mathbf{q}^{n+1}/\rho^{n+1}$. Panel IV, Optional POD compression of neural basis coefficients. $\Psi_{p}$, $\Psi_{q}$, and $\Psi_{c}$ are the respective reduced POD bases.
4.1 Residual–error certificate for the weighted Darcy projection